# Developing a Thailand solar irradiance map using Himawari-8 satellite imageries and deep learning models

### v2

@@ Abstract
This paper presents an online platform showing Thailand solar irradiance map every 30 minutes available at https://www.cusolarforecast.com. The methodology for estimating global horizontal irradiance (GHI) across Thailand relies on cloud index extracted from Himawari-8 satellite imagery, Ineichen clear-sky model with locally-tuned Linke turbidity, and machine learning models. The methods take clear-sky irradiance, cloud index, re-analyzed GHI and temperature data from the MERRA-2 database, and date-time as inputs for GHI estimation models, including LightGBM, LSTM, Informer, and Transformer. These are benchmarked with the estimate from a commercial service X by evaluation of 15-minute ground GHI data from 53 ground stations over 1.5 years during 2022-2023. The results show that the four models have competitive performances and outperform the service X. The best model is LightGBM with an MAE of 78.58 and RMSE of 118.97 . Obtaining re-analyzed MERRA-2 data for the whole Thailand region is not economically feasible for deployment. When removing these features, the Informer model has a winning performance in MAE of 78.67 . The obtained performance aligns with existing literature by taking the climate zone and time granularity of data into consideration. As the map shows an estimate of GHI over 93,000 grids with a frequent update, the paper also describes a computational framework for displaying the entire map. It tests the runtime performance of deep learning models in the GHI estimation process.
@@ 1.3 Neural network techniques
A methodology that exploits deep-learning or tree-based models to capture a complex relationship between satellite cloud and GHI. To find the best model, we benchmark LSTM, Informer, Transformer, and LightGBM (LGBM) for GHI estimation. Additionally, we improved Informer and Transformer particularly for this application. For LSTM and LGBM, we study the impact of input features. The results show that all these models provide more accurate GHI estimates than the one from a commercial service X. Our implementation are provided at https://github.com/energyCUEE/solarmap-DL-2024.
@@ 2.2 Data resources
There are 576 solar sites registered to the Energy Regulatory Commission of Thailand (ERC), where information on locations and installed capacity are provided. Ground measurements of GHI and satellite cloud data are required for model training and validation among these sites. Figure 1 (b), shows 53 sites that record solar irradiance, solar power, ambient and PV module temperature measurements, every 15 minutes from October 2020 to June 2023.
@@ 4.3 Satellite-derived GHI estimation
Figure 11 shows a comparative performance of all GHI estimation models, including LGBM, LSTM, Informer, Transformer, and the service X. All the proposed benchmarked models provide a similar MAEs, that are 20, 78, 98 in clear, partly cloudy, and cloudy sky conditions, respectively, except for the MAE of the service X that is the highest in all conditions; see Figure 11 (Top.). For the MBE, LGBM and the service X have a relatively high negative bias (underestimate) in clear and partly cloudy and with positive bias (overestimate) in the cloudy condition. Meanwhile, LSTM, Informer, and Transformer tend to overestimate—low positive bias in clear and partly cloudy and high positive bias in the cloudy condition, as shown in Figure 11 (Bottom.) The evaluation with the RMSE metric is reported in Table 3. All the proposed models perform almost equally if MAE is considered. However, considering RMSE performance, LGBM has the lowest error in all sky-condition settings. This is because the training loss of the tree model is MSE, while the training loss of deep learning models is set to MAE loss.
@@
| Overall MAE | 78.58 | 80.41 | 78.97 | 78.52 | 89.97 |
| Clear sky | 19.73 | 18.92 | 17.99 | 18.11 | 47.67 |
| Partly cloudy | 75.93 | 77.62 | 76.30 | 75.90 | 89.19 |
| Cloudy | 97.01 | 99.68 | 97.81 | 97.07 | 99.65 |
@@
| Overall RMSE | 118.97 | 125.86 | 124.54 | 123.82 | 129.53 |
| Clear sky | 36.70 | 32.79 | 31.54 | 31.94 | 58.41 |
| Partly cloudy | 115.13 | 122.19 | 121.21 | 120.58 | 127.44 |
| Cloudy | 138.53 | 145.74 | 143.45 | 142.42 | 144.05 |
@@
| Overall MBE | -0.47 | 6.11 | 4.55 | 5.59 | -4.73 |
| Clear sky | -7.86 | -1.24 | -1.86 | -1.71 | -7.67 |
| Partly cloudy | -7.48 | 2.04 | 0.53 | 2.09 | -6.74 |
| Cloudy | 23.07 | 20.29 | 18.41 | 17.97 | 2.17 |

### v3

@@ Abstract
This paper presents an online platform showing Thailand solar irradiance map every 30 minutes available at https://www.cusolarforecast.com. The methodology for estimating global horizontal irradiance (GHI) across Thailand relies on cloud index extracted from Himawari-8 satellite imagery, Ineichen clear-sky model with locally-tuned Linke turbidity, and machine learning models. The methods take clear-sky irradiance, cloud index, re-analyzed GHI and temperature data from the MERRA-2 database, and date-time as inputs for GHI estimation models, including LightGBM, LSTM, Informer, and Transformer. These are benchmarked with the estimate from a commercial service X by evaluation of 15-minute ground GHI data from 53 ground stations over 1.5 years during 2022-2023. The results show that the four models exhibits comparable overall MAE performance to the service X. The best model is LightGBM with an overall MAE of 78.58 and RMSE of 118.97 , while the service X achieves the lowest MAE, RMSE, and MBE in cloudy condition. Obtaining re-analyzed MERRA-2 data for the whole Thailand region is not economically feasible for deployment. When removing these features, the Informer model has a winning performance in MAE of 78.67 . The obtained performance aligns with existing literature by taking the climate zone and time granularity of data into consideration. As the map shows an estimate of GHI over 93,000 grids with a frequent update, the paper also describes a computational framework for displaying the entire map. It tests the runtime performance of deep learning models in the GHI estimation process.
@@ 1.3 Neural network techniques
A methodology that exploits deep-learning or tree-based models to capture a complex relationship between satellite cloud and GHI. To find the best model, we benchmark LSTM, Informer, Transformer, and LightGBM (LGBM) for GHI estimation. Additionally, we improved Informer and Transformer particularly for this application. For LSTM and LGBM, we study the impact of input features. Our findings indicate that all these models exhibit overall performance comparable to commercial service X, a widely recognized industry benchmark. Our implementation are provided at https://github.com/energyCUEE/solarmap-DL-2024.
@@ 2.2 Data resources
There are 576 solar sites registered to the Energy Regulatory Commission of Thailand (ERC), where information on locations and installed capacity are provided. Ground measurements of GHI and satellite cloud data are required for model training and validation among these sites. Figure 1 (b), shows 53 solar farms that record 15-minute average of solar irradiance, solar power, ambient and PV module temperature measurements from October 2020 to June 2023. The pyranometer model of solar irradiance measurement is SMP 6 (ISO 9060 spectrally flat class B). To ensure data quality, a pre-processing pipeline was implemented.
Initially, data was filtered to retain only daytime measurements. GHI values were restricted to a physically realistic range of 0 to 1400 . Dates with excessively low average daytime GHI were discarded. This step aimed to eliminate periods with insufficient solar radiation or inaccurate measured values. Dates with missing data periods exceeding six hours were entirely removed. Shorter data gaps were filled using linear interpolation. Last, a manual inspection of GHI data and associated variables for consistency was performed. As a result of this quality control process, three of the original 56 sites were unsuitable and excluded from the analysis. The remaining 53 sites were deemed sufficiently reliable for this study.
@@
The data during the period of satellite outage or error image scan are flagged to be imputed. If the imputation period is small, a simple interpolation from nearby data points is performed, while the missing data in a long period are removed from the study.
@@
Pre-processing steps were applied to cloud images and extracted cloud index (CI) data to ensure alignment with the target variable, GHI measurements. Initially, satellite images containing significant areas of bad scan (black area) were discarded. For each day, the first and last cloud image within the 7:00 to 17:00 timeframe were identified. Missing cloud images within this range were imputed using a forward-fill method. To accommodate the GHI estimation model’s requirement for five lagged CI values, a backward-fill approach was employed for the first five time steps. For the imputed date-time between 5:00 and 6:00, zero-filling was applied; otherwise, backward-filling from the first image’s CI value was utilized.
@@ 4.3 Satellite-derived GHI estimation
Figure 11 and Table 3 show a comparative performance of all GHI estimation models, including LGBM, LSTM, Informer, Transformer, and the service X. All the proposed benchmarked models provide similar overall MAEs in the range of 78-81 . For the MAE, the Informer, Transformer, and service X have the lowest error in clear-sky, partly cloudy, and cloudy conditions, respectively. For the MBE, LGBM and the service X have a relatively high negative bias (underestimate) in clear and partly cloudy condition. Meanwhile, LSTM, Informer, and Transformer tend to overestimate—high positive bias in the cloudy condition, while the service X has the lowest MBE in this case, as shown in Figure 11 (Bottom.) The evaluation with the RMSE metric is reported in Table 3. All the proposed models perform almost equally if MAE is considered. For specific sky conditions, our proposed models generally perform well in clear-sky and partly cloudy scenarios. However, service X demonstrates superior performance under cloudy conditions.
@@
| Overall MAE | 78.58 | 80.41 | 78.97 | 78.52 | 80.80 |
| Clear sky | 19.73 | 18.92 | 17.99 | 18.11 | 27.00 |
| Partly cloudy | 75.93 | 77.62 | 76.30 | 75.90 | 80.16 |
| Cloudy | 97.01 | 99.68 | 97.81 | 97.07 | 91.99 |
@@
| Overall RMSE | 118.97 | 125.86 | 124.54 | 123.82 | 121.99 |
| Clear sky | 36.70 | 32.79 | 31.54 | 31.94 | 39.52 |
| Partly cloudy | 115.13 | 122.19 | 121.21 | 120.58 | 120.58 |
| Cloudy | 138.53 | 145.74 | 143.45 | 142.42 | 135.01 |
@@
| Overall MBE | -0.47 | 6.11 | 4.55 | 5.59 | -4.21 |
| Clear sky | -7.86 | -1.24 | -1.86 | -1.71 | -5.62 |
| Partly cloudy | -7.48 | 2.04 | 0.53 | 2.09 | -6.17 |
| Cloudy | 23.07 | 20.29 | 18.41 | 17.97 | 2.26 |